Cj wants to mix cashew and almonds. Cashews cost 2 dollars per pound and almonds cost 5 dollars per pound. He plans to sell 150 pounds of a mixture. How many pounds of each type of nuts should be mixed if the mixture will cost 3 dollars?
Answer by fractalier(6550)

Let the amount of cashews be x. Then the amount of almonds will be 150-x.
The set up looks like this
2x + 5(150-x) = 3(150)
2x + 750 - 5x = 450
-3x + 750 = 450
-3x = -300
x = 100 pounds of cashews
150-x = 50 pounds of almonds